OpenStudy (anonymous):

Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of ten inches and a height of 3 inches. Each leg is 3 inches by 3 inches by 12 inches. How much area will Evan have to paint? Use three point one four for pi. Show your work.

OpenStudy (anonymous):

hii , I do not know how I defferatioate the following eqation with respect to ci and with respect to c0 \[\epsilon (c)=< c0\theta0+\sum_{i}ci \theta i|\nu|c0\theta 0+\sum_{i}ci \theta i >/c _{2}^{0}+g \sum_{i}c_{i}^{2}\]

OpenStudy (anonymous):

I will be very appreciated for any help